# Index Terms: Weakly supervised video anomaly detection, event prompt, multi-prompt learning, text.I Introduction

### v1

@@ IV Experiments
| ✓ | 85.2 | 64.6 | 93.7 | 80.5 | 88.7 | |||
| ✓ | ✓ | 87.0 | 68.6 | 94.2 | 81.8 | 91.3 | ||
| ✓ | ✓ | ✓ | 87.4 | 69.1 | 94.3 | 82.4 | 91.4 | |
| ✓ | ✓ | ✓ | ✓ | 87.7 | 71.1 | 94.4 | 82.6 | 93.1 |
IV-A Implementation Details
The dimension of the visual features extracted by CLIP [109] is 512, while the dimension of semantic features is 768. The prompt dictionary capacity is set to 30 for UCF-Crime, XD-Violence and TAD datasets, 25 for ShanghaiTech. The batch size is set to 64 for TAD dataset, and it is halved to 32 on the other three datasets. The number of snippets per video is set to 64 for all datasets. The hyper-parameters , and are consistent across all datasets. The Adam optimizer is utilized with a learning rate of 0.001 and weight decay of 0.005 during the training process.
IV-B Comparison Results
Quantitative analysis. The comparisons between our LAP model and other state-of-the-art (SOTA) WS-VAD models on the XD-Violence and UCF-Crime datasets are presented in Table I. It can be seen that the proposed model outperforms all the other methods in both datasets. Specifically, our model achieves the highest AP of 82.6% on the XD-Violence dataset, surpassing the second-highest method UR-DMU [103] by 0.9%. It is worth noting that, although TEVAD [104] also introduced semantic information by a text branch, its AP is 2.8% lower than ours. On the UCF-Crime dataset, the AUC of our LAP model (87.7%) is at least 0.8% higher than those most recent methods, such as UR-DMU [103] (87.0%), CMRL [107] (86.1%) and TEVAD [104] (84.9%). Compared to UMIL [101] (86.7%), which also uses a CLIP-based pre-trained model as its feature extractor, our model shows a noticeable improvement of 1.0%.
More comparisons on TAD and ShanghaiTech datasets are given in Table II. It shows that the AUC of our approach (93.1%) is still higher than all compared SOTA methods [101, 102, 106] by a margin of 0.2% to 3.9% on TAD dataset. Our method achieves the second highest AUC in ShanghaiTech, which is 0.2% lower than the best method CMRL [107]. Noting that, if the visual extractor is switched to I3D [93], the AUC will be boosted to 98.0% (0.4% higher CMRL [107]), which will be discussed in Section IV-D. In summary, our LAP consistently outperforms most compared SOTA methods on four challenging benchmarks of WS-VAD.
Qualitative analysis. To further demonstrate the effectiveness of our method, the qualitative comparisons between our approach (LAP) and the SOTA method TEVAD [104] are visualized in Figure 4.The normal and abnormal frames of videos from UCF-Crime and XD-Violence datasets are presented along with their corresponding frame-level anomaly scores, while green and red dashed rectangles indicate normal and abnormal ones, respectively. As shown in the figures, our method not only outperforms TEVAD [104] in terms of anomaly detection ability but also reduces false alarms on normal parts.
IV-C Ablation Studies
IV-C1 Components.
The proposed prompt-related components, i.e. feature synthesis (FS), multi-prompt learning (MPL) and pseudo anomaly labeling (PAL) are the keys to our superior performance in VAD. The ablation results of these three components on three datasets are shown in Table III. The baseline module is a visual-only branch MIL-based network with a CLIP feature extractor [109]. By cooperating text branch for feature synthesis, our model achieves 1.8%, 0.5% and 2.6% AUC improvement, respectively, which shows the efficiency of semantic information. The MPL module can also improve the AUC for all datasets by 0.4%, 0.1% and 0.1%, while the AP on XD-Violence and AUC on UCF-Crime are also boosted by 0.6% and 0.5%. Further incorporating the PAL module yields better results. It outperforms the baseline by 2.5% in AUC and 6.5% in AUC on UCF-Crime, 4.4% in AUC on TAD, as well as 0.7% in AUC and 2.1% in AP on XD-Violence.
IV-C2 Prompt format.
The prompt format plays an important role in the proposed LAP model. Thus, two different formats are tested in this experiment. One is organized by anomaly phrases, such as ”falling down” or ”on fire”. The other contains complete anomaly sentences, like ”someone is doing something to whom” or ”something is what”. As shown in Figure LABEL:fig:a, the sentence-based prompt dictionary outperforms the phrase-based one by 0.3% on XD-Violence and 0.4% on UCF-Crime, respectively. It suggests that prompts containing richer information are more helpful in identifying suspected anomalies.
IV-C3 Pseudo anomaly threshold.
As required by the PAL module, pseudo labels are determined according to the threshold . The dynamic threshold given in Eq. 13 is used in previous experiments, which is based on data distribution in the current batch. Another static threshold strategy is compared in this test, while is set to 0.5 as prior knowledge. As shown in Figure LABEL:fig:b, the dynamic threshold is better than the static one, whose AP and AUC are 0.1% and 0.4% higher on XD-Violence and UCF-Crime datasets.
IV-D Discussions
IV-D1 Class-wise AUC.
To demonstrate the detailed performance on specific abnormal events, the class AUC of our model is compared with RTFM [13] in Figure 6. It shows that the proposed LAP model outperforms RTFM in most categories, especially on “Arson”, “Assault”, “Explosion”, “RoadAccidents” and “Robbery”. This can be attributed to the effective use of our prompt dictionary to describe those anomalies including representative texts such as “fire”, “knife” or “accident”. Combined with the MPL module, their synthetic features are more likely to be identified as abnormal ones. However, our model may be less effective in some cases if the action is subtle or difficult to describe, such as “Shoplifting” and “Fighting” in Figure 6.
| Source | UCF-Crime | XD-Violence | XD-Violence | UCF-Crime | ST | UCF-Crime | XD-Violence |
| Target | UCF-Crime (AUC %) | XD-Violence (AP %) | ST (AUC %) | ||||
| RTFM [13] | 84.3 | 68.6 (-15.7) | 76.6 | 37.3 (-39.3) | 97.2 | 45.5 (-51.7) | 40.2 (-57.0) |
| CMRL [107] | 86.1 | 69.9 (-16.2) | 81.3 | 46.7 (-34.6) | 97.6 | 52.0 (-45.6) | 38.0 (-59.6) |
| Ours | 87.7 | 82.1 (-5.6) | 82.6 | 68.1 (-14.5) | 97.4 | 69.7(-27.7) | 87.5(-9.9) |
@@ IV-D Discussions
| Ours | 87.7 | 87.1(-0.6) | 85.7(-2.0) | 87.0(-0.7) |
IV-D2 Open set VAD.
@@
In practical applications, it is impossible to collect or define all possible anomalies in advance. Hence, it is crucial to examine the robustness of anomaly detection models when confronted with open abnormal categories in real-world scenarios. Following the protocol of open set VAD in MLAD [22], experiments are conducted on the top 3 largest anomaly categories from the UCF-Crime dataset, namely “Explosion”, “RoadAccidents” and “Shoplifting”. These categories are sequentially removed from the training set and treated as real open abnormal events. The comparisons with three SOTA models are presented in Table V. It is obvious that the proposed LAP model outperforms RTFM [13], MLAD [22] and TEVAD [104] in all three categories. It is worth noting that our method achieves minimal decreases in AUC values when compared to alternative approaches. This indicates that our method is more efficient in handling open abnormal event issues.
IV-D3 Cross-dataset performance.
@@
The definition of anomalies varies from different VAD datasets. For instance, the abnormal events in ShanghaiTech dataset (such as “walking across the road” and “skating”) are campus activities, which is quite different from the abnormal categories found in social datasets UCF-Crime and XD-Violence (such as “shooting” and “firing”). Thus, it will become a challenging transfer learning task, if the model is trained and inferred on different datasets. However, it is actually what will happen in real-world anomaly detection applications. To evaluate the generalization ability of our proposed method, another set of experiments using different sources of training and inference videos is conducted. Compared with RTFM [13] and CMRL [107], our model explains the definition of anomalies with their descriptions using the prompt dictionary and multi-prompt learning scheme. Such a new paradigm of utilizing the semantic information leads to an extraordinary cross-dataset performance as shown in Table IV. The performance degradation of the proposed model is only one-sixth of the ones of RTFM and CMRL, when it is trained on XD-Violence and tested on ShanghaiTech. It indicates that our method is much less sensitive to variations in the data domain, which is important for practical applications.
| Year | Method | Feature | AUC(%) |
| 2020 | AR-Net [108] | I3D | 91.2 |
| AR-Net+MPL+PAL | I3D | 92.5(+1.3) | |
| 2021 | RTFM [13] | I3D | 97.2 |
| RTFM+MPL+PAL | I3D | 98.0 (+0.8) | |
| 2022 | MCR [62] | I3D | 94.8 |
| MCR+MPL+PAL | I3D | 95.2 (+0.4) | |
| 2023 | CMRL [107] | I3D | 97.6 |
| UMIL [101] | X-CLIP | 96.8 |
IV-D4 Plug and play.
@@
To further explore the potential of our method, the proposed MPL and PAL modules are embedded into three representative WS-VAD works (AR-Net [108], MCR [62] and RTFM [13]) on ShanghaiTech. For a fair comparison, I3D [93], instead of CLIP [109], is used as the feature extractor, while all experimental settings were kept the same as in their respective papers. As shown in Table VI, the reimplemented frameworks generally exhibited better performance. All three LAP embedded frameworks achieve higher AUC than their vanilla forms with noticeable improvements (1.3%, 0.5%, 0.4%) on the ShanghaiTech dataset. Moreover, the re-implemented RTFM model could even surpass the latest SOTA model CMRL [107] by 0.4%. It indicates that more frameworks may benefit from our prompt-related modules, which are plug-and-play.

### v2

@@ IV-D Ablation Studies
| ✓ | 87.0 | 67.0 | 93.2 | 81.3 | 93.7 | |||
| ✓ | ✓ | 88.2 | 70.4 | 95.0 | 84.1 | 94.1 | ||
| ✓ | ✓ | ✓ | 88.5 | 70.7 | 95.4 | 85.0 | 94.2 | |
| ✓ | ✓ | ✓ | ✓ | 88.9 | 73.0 | 95.6 | 86.5 | 94.4 |
@@
| Ours | 88.9 | 88.1(-0.8) | 87.0(-1.9) | 88.4(-0.5) |
IV-E Discussions
Class-wise AUC. To demonstrate the detailed performance on specific abnormal events, the class AUC of our model is compared with RTFM [11] in Figure 7. It shows that the proposed LAP model outperforms RTFM in most categories, especially on “Assault”, “Explosion”, “RoadAccidents” and “Robbery”. This can be attributed to the effective use of our prompt dictionary to describe those anomalies including representative texts such as “fire”, “knife” or “accident”. Combined with the MPL module, their synthetic features are more likely to be identified as abnormal ones. However, our model may be less effective in some cases if the action is subtle or difficult to describe, such as “Shoplifting” and “Fighting” in Figure 7.
@@ IV-E Discussions
Open set VAD. In practical applications, it is impossible to collect or define all possible anomalies in advance. Hence, it is crucial to examine the robustness of anomaly detection models when confronted with open abnormal categories in real-world scenarios. Following the protocol of open set VAD in MLAD [47], experiments are conducted on the top 3 largest anomaly categories from the UCF-Crime dataset, namely “Explosion”, “RoadAccidents” and “Shoplifting”. These categories are sequentially removed from the training set and treated as real open abnormal events. The comparisons with three SOTA models are presented in Table IV. It is obvious that the proposed LAP model outperforms RTFM [11], MLAD [47] and TEVAD [12] in all three categories. It is worth noting that our method achieves minimal decreases in AUC values when compared to alternative approaches. This indicates that our method is more efficient in handling open abnormal event issues.
| Source | UCF | XD | XD | UCF |
| Target | UCF (AUC %) | XD (AP %) | ||
| RTFM [11] | 84.3 | 68.6 (-15.7) | 76.6 | 37.3 (-39.3) |
| CMRL [36] | 86.1 | 69.9 (-16.2) | 81.3 | 46.7 (-34.6) |
| Ours | 88.9 | 83.5 (-5.4) | 86.5 | 60.9(-25.6) |
@@
Cross-dataset performance. The categories of anomalies varies from different VAD datasets. For instance, the abnormal events in UCF-Crime dataset are collected from surveillance videos, which is quite different from the abnormal categories in XD-Violence developed from movies and online videos. Thus, it will become a challenging transfer learning task, if the model is trained and inferred on different datasets. However, it is actually what will happen in real-world anomaly detection applications. To evaluate the generalization and zero-shot abilities of our proposed method, another set of experiments using different sources of training and inference videos is conducted. Compared with RTFM [11] and CMRL [36], our model explains the definition of anomalies with their descriptions using the prompt dictionary and multi-prompt learning scheme. Such a new paradigm of utilizing the semantic information leads to an extraordinary cross-dataset performance as shown in Table V. The performance degradation of the proposed model is only one-third of the ones of RTFM and CMRL, when it is trained on XD-Violence and tested on ShanghaiTech. It indicates that our method is much less sensitive to variations in the data domain, which is important for practical applications.
| Method (ST) | AUCall | AUCabn | FARall | FARabn |
| RTFM | 97.2 | 64.3 | 0.06 | 0.86 |
| RTFM+MPL | 97.6 | 72.2 | 0.06 | 0.71 |
| RTFM+PAL | 97.5 | 73.9 | 0.03 | 0.44 |
| RTFM+Both | 98.0 | 75.6 | 0.04 | 0.58 |
@@
Plug and play. To further explore the potential of our method, the proposed MPL and PAL modules are embedded into the representative WS-VAD work RTFM [11] on ShanghaiTech. For a fair comparison, I3D [8], instead of CLIP [15], is used as the feature extractor, while all experimental settings were kept the same as in RTFM paper. As shown in Table VI, the reimplemented frameworks generally exhibited better performance. By incorporating either MPL or PAL alone, enhancements of 0.4% and 0.3% can be achieved on AUCall, whereas more significant enhancements of 7.9% and 9.6% can be observed on AUCabn. The efficacy of MPL and PAL is demonstrated by their ability to improve the performance of the conventional WSAD framework. Through the collaboration of MPL and PAL, LAP integrated RTFM demonstrates superior AUC and reduced FAR compared to its original version, showing significant enhancements (0.8%, 0.02%) on the ShanghaiTech dataset. It is worth noting that the reimplemented RTFM model (98. 0%) could even surpass the latest SOTA model CMRL [36] by 0.4%. It indicates that more frameworks may benefit from our prompt-related modules, which are plug-and-play.